A discount store sold plastic cups for $3.25 each and ceramic cups for $4.50 each. If 500 cups were sold for a total of $2090, how many cups of each type were sold?

Answer by mananth(16946)

A discount store sold plastic cups for $3.25 each and ceramic cups for $4.50 each. If 500 cups were sold for a total of $2090, how many cups of each type were sold?
...
plastic cups for $3.25 each ----------p numbers
ceramic cups for $4.50 each.----------c numbers
500 cups were sold
for a total of $2090
..
p+c=500......................(1)
3.25p+4.5c=2090...............(2)
..

multiply (1) by -3.25
we get
-3.25p-3.25c=-1625
add to (2)
..
3.25p-3.25p+4.5c-3.25c=2090-1625
1.25c=465
/1.25
c=372.
ceramic cups 372
Balance 128 plastic cups
...
m.ananth@hotmail.ca